OpenStudy (anonymous):

What is the solution to the expression ab − c, if a = 5, b = 3, and c = 2? 10 12 13 15

jhonyy9 (jhonyy9):

13

OpenStudy (anonymous):

do you have an explanation

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

i got it

jhonyy9 (jhonyy9):

5*3=15-2=13

jhonyy9 (jhonyy9):

medal ?

jhonyy9 (jhonyy9):

thank you